# Generalized Bayesian Multidimensional Scaling and Model Comparison

### v2

@@ 6 Conclusion and Discussion
In this work, we introduced a generalized Bayesian multidimensional scaling (GBMDS) framework that addresses the limitations of existing BMDS methods based on MCMC. By incorporating non-Gaussian error distributions and diverse dissimilarity metrics, we enhanced the robustness and flexibility of MDS. We also developed an adaptive ASMC algorithm for efficient Bayesian inference, which provides a nearly unbiased marginal likelihood estimator as a byproduct. This estimator facilitates model comparison across candidate models with different error distributions, dissimilarity metrics, and dimensions. Our approach was validated using both synthetic and real data, demonstrating its superior computational efficiency, scalability, and robustness compared to traditional MCMC-based methods under the same computational budget.
@@
Bayesian model selection offers several advantages over classical approaches such as $p$-values and information criteria. First, Bayes factors provide a direct probabilistic measure of model support, making them intuitive even for non-statisticians. Second, Bayesian model selection is consistent, meaning that given enough data, it will asymptotically select the model with the smallest Kullback-Leibler divergence to the true model under mild conditions (Berk, 1966), whereas some classical methods, such as hypothesis testing based on $p$-values, do not always guarantee consistency. Third, Bayesian model selection naturally penalizes excessive model complexity, favoring simpler models when competing models fit the data similarly (Berger et al., 2001; Robert et al., 2007).
@@
Due to these advantages, we propose selecting the optimal model-dimension combination for GBMDS using Bayes factors. Unlike the simple Bayesian criterion in Oh and Raftery (2001), our fully Bayesian approach directly utilizes the nearly unbiased marginal likelihood estimator from the annealed SMC algorithm with an adaptive annealing scheme, providing a principled and computationally efficient framework for the choice of model and dimension. In contrast, frequentist model selection via STRESS relies on the specifically constructed statistics that are tailored to particular cases, limiting its generalizability. Our approach estimates the marginal likelihood as a byproduct of annealed SMC, eliminating the need for additional computational steps. This efficient estimation procedure not only facilitates robust model comparison across different distance metrics and dimensional choices but also provides a significant computational advantage over existing MCMC-based methods.
The proposed annealed SMC algorithm offers a viable alternative to MCMC, particularly when computational resources are available and faster inference is prioritized over long MCMC chains. Note that the proposed adaptive ASMC framework is highly flexible and can straightforwardly integrate existing MCMC kernels from the literature, including those based on HMC, to further enhance computational efficiency. While memory constraints can arise for high-dimensional particles, dimension reduction methods such as MDS yield low-dimensional latent variables, making memory demands more manageable. Moreover, our method does not require storing the complete history of particles at each annealing step, further reducing memory overhead. As a result, even a personal laptop can efficiently handle a large number of particles. Memory constraints can also be addressed by either increasing the number of SMC iterations while using fewer particles—an approach that has been shown to maintain or even improve performance with proper tuning (Wang et al., 2020)—or by leveraging parallel implementations that distribute the particle set across computing cores. Methods such as entangled Monte Carlo (Jun et al., 2012), the island particle filter (Vergé et al., 2015), and the distributed particle filter (Heine and Whiteley, 2017) enhance scalability, ensuring that memory limitations do not pose a significant barrier to running SMC with a large number of particles.
The proposed method is implemented in an R package GBMDS which can be accessed at https://github.com/SFU-Stat-ML/GBMDS. The implementation integrates R and C++ to enhance computational efficiency, making it more accessible for practitioners. To assess scalability and runtime performance, we compared our implementation with the MCMC-based method in Oh and Raftery (2001) using their available R package BayMDS (Oh and Lee, 2022) across various simulation studies and data applications. The results indicate that even for more complex models, such as the truncated skewed Gaussian and truncated Student’s $t$, our method achieves comparable or faster runtimes, demonstrating its scalability to larger datasets. When applying the BayMDS package to datasets with over 1,000 observations, we encountered processing limitations due to the requested data size exceeding its computational capacity. This further highlights the practical advantages of our proposed approach in handling large-scale problems.
We have applied Procrustes transformations to the posterior samples to address the non-identifiability issue. The transformed posterior samples were then used to construct credible regions, providing a measure of uncertainty. In the geographical data example, we observed that the credible regions for certain observations were relatively broad, suggesting that specific patterns should be interpreted with caution. Exploring alternative post-processing transformations could further enhance interpretability. Additionally, it would be valuable to investigate how the number of particles in the annealed SMC algorithm affects the credible regions.
There are several directions for future work. Our proposed framework aids in identifying a suitable model for dissimilarities and an optimal $p$ for dimension reduction. While Euclidean models are well-suited for two-dimensional visualization, representing object configurations in a non-Euclidean space remains an open challenge. Future work could explore effective visualization techniques for such structures, improving interpretability in high-dimensional non-Euclidean embeddings. Another important application of MDS is clustering, where observations are grouped based on dissimilarity. In our final data application, we evaluated the proposed model using heuristic clustering methods. However, model-based clustering provides a probabilistic alternative that allows for more rigorous classification and uncertainty quantification. Building on Oh and Raftery (2007), one could extend the adaptive annealed SMC algorithm to jointly perform MDS and model-based clustering, providing a fully Bayesian treatment of clustering assignments. This extension could also apply to text clustering, where carefully chosen dissimilarity metrics define meaningful cluster structures. Additionally, future work could focus on enhancing the efficiency of the annealed SMC algorithm. One straightforward approach is to integrate existing MCMC kernels from the BMDS literature into ASMC for comparison. Another promising direction is subsampling-based inference (Gunawan et al., 2020), which can reduce computational costs while maintaining accuracy. Furthermore, advanced parallelization strategies could significantly improve scalability, enabling the proposed method to accommodate increasingly large datasets.
@@
[…]

### v3

@@ 6 Conclusion and Discussion
In this work, we introduced a generalized Bayesian multidimensional scaling (GBMDS) framework that addresses the limitations of existing MCMC-based BMDS methods. By incorporating non-Gaussian error distributions and diverse dissimilarity metrics, we enhanced the robustness and flexibility of MDS. We also developed an adaptive ASMC algorithm for efficient Bayesian inference. Unlike standard MCMC, which requires separate, often complex procedures to estimate model evidence, our algorithm provides a marginal likelihood estimator naturally as a byproduct of sampling. This facilitates rigorous model comparison across candidate models with varying error distributions, dissimilarity metrics, and dimensions.
@@
Bayesian model selection offers distinct advantages over classical approaches such as $p$-values and information criteria. First, Bayes factors provide a direct probabilistic measure of model support, offering intuitive interpretation. Second, Bayesian model selection is consistent, meaning that given sufficient data, it asymptotically selects the model minimizing the Kullback-Leibler divergence to the true model under mild conditions [berk1966limiting]. Third, it naturally penalizes excessive model complexity, favoring parsimony[berger2001objective, robert2007bayesian].
@@
Leveraging these advantages, we propose selecting the optimal model-dimension combination for GBMDS using Bayes factors. While frequentist model selection via STRESS relies on loss statistics tailored to particular cases, our fully Bayesian approach utilizes the marginal likelihood estimates from the adaptive ASMC algorithm. Although adaptive schemes may introduce negligible bias, they provide significant gains in robustness and variance reduction compared to fixed schedules. This framework eliminates the need for additional computational steps, offering a principled and efficient alternative to existing MCMC-based estimators.
The proposed annealed SMC algorithm offers a robust alternative to MCMC, particularly when computational resources permit parallelization or when robust inference is prioritized over single-chain MCMC exploration. The adaptive ASMC framework is highly flexible, allowing for the straightforward integration of advanced MCMC kernels from existing literature, such as Hamiltonian Monte Carlo (HMC), to further enhance mixing efficiency. Regarding memory constraints, while high-dimensional particle systems can present challenges, these are mitigated in BMDS by the typically low dimensionality of the latent embeddings. Furthermore, our method eliminates the need to store the complete particle history at every annealing step, significantly reducing overhead; consequently, even standard personal computing hardware can efficiently handle substantial particle populations. Should memory limitations arise in large-scale applications, they can be effectively managed by trading particle count for annealing iterations, which is a strategy shown to maintain or improve performance with appropriate tuning [wang2020annealed]. Alternatively, scalability can be ensured through distributed computing strategies such as entangled Monte Carlo [NIPS2012_dc4c44f6], the island particle filter [Verge_island_smc_2015], and the distributed particle filter [HEINE20172508], ensuring that memory limitations do not hinder the application of SMC to large datasets
The proposed method is implemented in the R package GBMDS (https://github.com/SFU-Stat-ML/GBMDS), integrating R and C++ for efficiency. Comparison with the MCMC-based package BayMDS [oh2022baymds] demonstrated that our method achieves comparable or faster runtimes even for complex models (e.g., truncated Student’s $t$), and successfully handles datasets exceeding 1,000 observations where traditional MCMC implementations faced capacity limitations.
Despite the methodological advancements of the proposed GBMDS framework, several limitations remain. First, addressing the inherent non-identifiability of the latent configuration relies on post-hoc Procrustes transformations. While this alignment allows for the construction of credible regions, we observed that these regions can remain relatively broad in scenarios with high noise or weak signal, potentially obscuring fine-grained spatial patterns. Second, although our adaptive algorithm significantly improves computational efficiency over MCMC, scalability to massive datasets remains a challenge. Specifically, the memory required to store and propagate the particle population grows linearly with both the sample size and the latent dimension. While our batch processing strategy mitigates this, extremely large datasets may still necessitate further approximation techniques or distributed computing implementations to remain computationally feasible. Third, the proposed adaptive algorithm currently assumes a fixed global latent dimension $p$ across all data batches. This assumption may be restrictive in streaming or sequential settings where the intrinsic dimensionality of the data varies over time or across partitions, necessitating more complex alignment or dimension-jumping mechanisms.
There are several directions for future work. First, while we utilized Procrustes transformations to address non-identifiability and visualize uncertainty, exploring alternative post-processing transformations could further enhance interpretability. Second, our current framework assumes a fixed latent dimension; future work should address cases where intrinsic dimensionality varies across data batches or partitions, potentially requiring alignment steps for valid comparison. Third, extending the adaptive ASMC algorithm to perform joint MDS and clustering could further enhance the versatility of the framework in detecting structure. Finally, future research should focus on enhancing the computational efficiency of the algorithm. Key directions include integrating established MCMC kernels from the recent BMDS literature, adopting subsampling-based inference [gunawan2020subsampling] to reduce computational costs, and implementing advanced parallelization strategies to improve scalability for increasingly large datasets.
@@
Supplementary information
Supplementary materials include a document with the details of the dissimilarity metrics, details of the particle propagation step, and details of the data-generating process in simulation studies.